OpenStudy (anonymous):

Write the equation of the line that passes by (–1, 4) with a slope of –1 in standard form

OpenStudy (anonymous):

slope = gradient = m equation of the line is of the form y = mx + c input known data 4 = -1 * -1 + c c = 3 equation is y = -x + 3